18hcenter:

Calculate the wavelength of a photon of energy 5.518 × 10-26 joules. (Planck's constant is 6.626 × 10-34 joule seconds, the speed of light is 2.998 × 108 m/s) 3.6 m 0.277 m

sillybilly123:

\(\color{red}{E_{\gamma} = h \nu }= h { c \over \lambda} \implies \lambda = {hc\over E_{\gamma}} = {6.626 \times 10^{-34} \times 2.998 \times 10^8 \over 5.518 \times 10^{-26}} = \dots \dots \)